The mean height of males 20 years or older is 68 inches with a standard deviation of 3.5 inches The mean height of females 20 years or older is 62 inches with a standard deviation of 2.5 inches. While a male has a height z-score of-1.3525, a female has a height z-score of 1.3525. How tall is he in inches? How tall is she in inches? 25. 26. The mean weight of males 20 years or older is 1701bs. with a standard deviation of 19 lbs. The mean weight of females 20 years or older is 145 lbs. with a standard deviation of 17 lbs. While a male is 180 lbs, a female is 154 1bs. What is his standardized weight (the z-score)? 27. 28 What is her standardized weight (the z-score) Test ifthe dataset {7,22,1 6,5,46,40,21,22,33,90} is normally distributed by using Shapiro-Wilk 29. What is the P-value? test. 30. Is it normally distributed? (Yes or No)
Answer #1

Solution25:

for males z=-1.3525

mean=68

sd=3.5

z=x-mean/sd

-1.3525=x-68/3.5

x=-1.3525*3.5+68

x=63.26625

x=63.3

he is 63.3 nches tall

Solution24:

For females mean=62

sd=2.5

z=1.3525

z=x-mean/sd

1.3525=x-62/2.5

x=1.3525*2.5+62

x=65.38125

x=65.4

she is 65.4 inches tall

Solution27:

z males=x-mean/sd

z=180-170/19

z=10/19

z=0.5263158

Solution28:

Z for females

z=x-mean/sd

=154-145/17

z=9/17

z=0.5294118